# Accelerated Parallel Tempering via Neural Transports

### v3

@@ Abstract
Markov Chain Monte Carlo (MCMC) algorithms are essential tools in computational statistics for sampling from unnormalised probability distributions, but can be fragile when targeting high-dimensional, multimodal, or complex target distributions. Parallel Tempering (PT) enhances MCMC’s sample efficiency through annealing and parallel computation, propagating samples from tractable reference distributions to intractable targets via state swapping across interpolating distributions. The effectiveness of PT is limited by the often minimal overlap between adjacent distributions in challenging problems, which requires increasing the computational resources to compensate. We introduce a framework that accelerates PT by leveraging neural samplers—including normalising flows, diffusion models, and controlled diffusions—to reduce the required overlap. Our approach utilises neural samplers in parallel, circumventing the computational burden of neural samplers while preserving the asymptotic consistency of classical PT. We demonstrate theoretically and empirically on a variety of multimodal sampling problems that our method improves sample quality, reduces the computational cost compared to classical PT, and enables efficient free energy/normalising constant estimation.
@@ 1 Introduction
Sampling from a probability distribution $\pi(x)=\exp(-U(x))/Z$ defined over a state-space $\mathcal{X}$ with tractable potential $U:\mathcal{X}\to\mathbb{R}$ and intractable normalising constant $Z=\int_{\mathcal{X}}\exp(-U(x))\mathrm{d}x$ is a fundamental task in machine learning and natural sciences. Markov Chain Monte Carlo (MCMC) methods are usually employed for such purposes, constructing an ergodic Markov chain $(X_{t})_{t\in\mathbb{N}}$ using local moves leaving $\pi$ invariant. While MCMC algorithms are guaranteed to converge asymptotically, in practice, they struggle when the target is complex with multiple well-separated modes (Papamarkou et al., 2022; Hénin et al., 2022). To handle such cases, Parallel Tempering (PT) (Swendsen & Wang, 1986; Geyer, 1991; Hukushima & Nemoto, 1996) is a popular class of MCMC methods designed to improve the global mixing of locally efficient MCMC algorithms.
@@
PT works by considering an annealing path $\pi^{0},\pi^{1},\dots,\pi^{N}$ of distributions over $\mathcal{X}$ interpolating between a simple reference distribution $\pi^{0}=\eta$ (e.g. a Gaussian) and the target $\pi^{N}=\pi$. PT algorithms construct a Markov chain $\textbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$, targeting the joint distribution $\bm{\pi}=\pi^{0}\otimes\cdots\otimes\pi^{N}$. The PT chain $\mathbf{X}_{t}$ is constructed by alternating between (1) a local exploration phase where the $n$-th chain  [^1] Following the PT literature, we also refer to components of $\mathbf{X}_{t}$ as chains. of $\mathbf{X}_{t}$ is updated according to a $\pi^{n}$-invariant MCMC algorithm; and (2) a communication phase which proposes a sequence of swaps between neighbouring states accepted according to a Metropolis–Hastings correction ensuring invariance (see Figure˜1 (a)). Crucially, PT offsets the additional computation burden of simulating the extended $N$ chains through parallel computation, allowing for a similar effective computational cost as a single chain when implemented in a maximally parallelised manner.
@@
Typically, the chains $X_{t}^{n}$ mix faster when closer to the reference and struggle closer towards the target. Therefore, communication between the reference and target, facilitated through swaps, can induce rapid mixing between modes of the target component (Woodard et al., 2009; Surjanovic et al., 2024). While the importance of the swapping mechanism for PT has led to a literature dedicated to optimising communication between the reference and target (Syed et al., 2021; 2022; Surjanovic et al., 2022) such works still rely on the original swapping mechanism (Geyer, 1991).
@@
Figure 1: (Left) An illustration of the local exploration and communication step for PT vs APT. (Middle) 1,000 samples of a Gaussian mixture model target obtained using PT vs APT with a standard Gaussian reference. See Section˜6.1 for more details. (Right) Round trips for PT and APT with $N=6$ chains over $T=100,000$ iterations of Algorithm˜1.
@@
We answer positively this question by formalising and exploiting the framework introduced by Ballard & Jarzynski (2009; 2012) in physics for designing more flexible swap mechanisms, which we call accelerated PT (APT). APT preserves PT’s asymptotic consistency and allows us to integrate easily normalising flows, stochastic control, and diffusions into existing PT implementations. ATP uses neural sampling in a parallelised manner, mitigating its high computational burden. Empirically, it outperforms PT by accelerating the communication between the reference and target states, even when controlling for the additional computation incurred by neural samplers.
A relevant prior work integrating PT with normalising flows is Invernizzi et al. (2022): it trains a normalising flow to directly map configurations from the highest-temperature reference distribution of a molecular system to the lowest-temperature target distribution, effectively bypassing the intermediate annealing distributions. By contrast, our framework leverages normalising flows to facilitate exchanges between all neighbouring temperature levels, thereby enhancing sampling efficiency across the entire annealing path and providing a more stable training objective.
@@ 2 Parallel Tempering
Let $\pi^{0},\pi^{1},\dots,\pi^{N}$ be an annealing path of probability distributions on $\mathcal{X}$ where $\pi^{0}=\eta$ is the reference and $\pi^{N}=\pi$ is the target. We assume the $n$-th annealing distribution admits density $\pi^{n}(x):=\exp(-U^{n}(x))/Z_{n}$ with respect to a base measure $\mathrm{d}x$, where $U^{n}:\mathcal{X}\to\mathbb{R}$ is the potential and $Z_{n}:=\int_{\mathcal{X}}\exp(-U(x))\mathrm{d}x$. We further assume we can evaluate the work $W^{n}:\mathcal{X}\to\mathbb{R}$ between $\pi^{n-1}$ and $\pi^{n}$, equal to the change in log-likelihood of the un-normalised annealing distributions:
$W^{n}(x):=\Delta F_{n}-\log\frac{\mathrm{d}\pi^{n}}{\mathrm{d}\pi^{n-1}}(x)=U^{n}(x)-U^{n-1}(x),$ (1)
where $\Delta F_{n}:=\log Z_{n-1}-\log Z_{n}$. Our goal is to estimate $\pi[f]:=\int_{\mathcal{X}}f(x)\pi(\mathrm{d}x)$, the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi$, and the change in free energy $\Delta F:=\log Z_{0}-\log Z$.
There is considerable flexibility in choosing annealing distributions provided $U^{0}(x)=\log\eta(x)$ with $Z_{0}=1$ and $U^{N}(x)=U(x)$ with $Z_{N}=Z$. Without loss of generality, we can assume $U^{n}=U^{\beta_{n}}$ where for $\beta\in[0,1]$, $U^{\beta}:\mathcal{X}\to\mathbb{R}$ continuously interpolates between reference and target potentials as $\beta$ increases from 0 to $1$, according to the annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$. A common choice is the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. See Masrani et al. (2021); Syed et al. (2021); Máté & Fleuret (2023); York (2023) for alternative non-linear annealing paths.
@@ 2.1 Non-Reversible Parallel Tempering
The PT algorithm constructs a Markov chain $\mathbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$ invariant to the joint distribution $\bm{\pi}:=\pi^{0}\otimes\cdots\otimes\pi^{N}$. We construct $\mathbf{X}_{t}$ from $\mathbf{X}_{t-1}$ by doing (1) a local exploration step followed by (2) a communication step seen in the top of Figure˜1 (a). For $n=0,\dots,N$, the $n$-th local exploration move (1) updates the $n$-th component of $\mathbf{X}_{t-1}$ using a $\pi^{n}$-invariant Markov kernel $K^{n}(x,\mathrm{d}x^{\prime})$ on $\mathcal{X}$ corresponding to an MCMC move targetting $\pi^{n}$,
@@
$\alpha^{n}(x,x^{\prime}):=\exp\left(\min\left\{0,W^{n}(x^{\prime})-W^{n}(x)\right\}\right).$ (2)
In practice, it is advantageous to use a non-reversible communication (Okabe et al., 2001; Syed et al., 2022), where the $n$-th swap move is proposed only at iterations with matching parity, $n\equiv t\mod 2$. Both local exploration and communication steps can be done in parallel, allowing distributed implementations to leverage parallel computation to accelerate sampling (Surjanovic et al., 2023).
Round trips
While the effective sample size (ESS) of samples generated by a Markov chain is the gold standard for evaluating the performance of MCMC algorithms, in our setting, we are mainly interested in improving the swap kernel within PT. As ESS measures the intertwined performance of the local exploration and swap kernels, we are instead interested in maximising communication between reference and target, which can be empirically measured by counting the total number of round trips, $R_{T}$ tracking the number of independent reference samples transported to the target after $T$ iterations of PT (Katzgraber et al., 2006; Lingenheil et al., 2009). See Section˜A.1 for a formal definition. The mixing time of the PT chain is related to the time it takes for a round trip to occur and the total number of round trips is a measure of the particle diversity generated by PT and strongly correlates with the ESS (Surjanovic et al., 2024).
@@
12:   if $\log U<\accentset{\rightharpoonup}{W}^{n}_{K,t}-\accentset{\leftharpoonup}{W}^{n}_{K,t}$ then $\triangleright$ Accelerated swap move
13:     $X^{n-1}_{t},X^{n}_{t}\leftarrow X^{n}_{t,0},X^{n-1}_{t,K}$
@@ 3 Accelerated Parallel Tempering
A limitation of PT is its inflexible swap move, which only proposes directly exchangeable samples between distributions, considering just the relative change in potentials $\pi^{n-1}$ and $\pi^{n}$. Low acceptance probability occurs when these distributions have minimal overlap. Addressing this requires increasing the number of parallel chains $N$, which may not always be possible. We propose Accelerated PT (APT), expanding the framework developed in Ballard & Jarzynski (2009; 2012) to improve distributional overlap and accelerate communication for a fixed number of chains.
@@ 3.1 Forward and Backward Accelerators
For $n=1,\dots,N$, let $P_{k}^{n-1}(x_{k-1},\mathrm{d}x_{k})$ and $Q_{k-1}^{n}(x_{k},\mathrm{d}x_{k-1})$, $k=1,\dots,K$, be two families of $K$ transition kernels which we call forward accelerators and backward accelerators. They induce ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$, two time-inhomogeneous Markov processes obtained by propagating $\pi^{n-1}$ forward in time and $\pi^{n}$ backward in time respectively using the forward and the backward accelerators,
@@
We assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ are mutually absolutely continuous and we can evaluate $W^{n}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$, the work between the forward and the backward accelerated paths, extending $W^{n}$ in Equation˜1,
$W^{n}_{K}(x_{0:K}):=\Delta F_{n}-\log\frac{\mathrm{d}{{\mathbb{Q}}}^{n}_{K}}{\mathrm{d}{{\mathbb{P}}}^{n-1}_{K}}(x_{0:K}).$ (3)
@@ 3.2 Non-Reversible Accelerated Parallel Tempering
We construct the Markov chain $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ for $t=1,\dots,T$ using the same local exploration and non-reversible communication as classical PT, but we use using the accelerated PT swap move as described below and summarized in Algorithm˜1. Given PT state $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})\in\mathcal{X}^{N+1}$ after the local exploration move, we define the $n$-th accelerated swap move as follows: generate paths $\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K}$, and $\accentset{\leftharpoonup}{X}^{n}_{t,0:K}$ obtained by propagating $X^{n-1}_{t}$ and $X^{n}_{t}$ forward and backward in time using the forward and backward transitions respectively,
@@
$\alpha_{K}^{n}(x_{0:K},x_{0:K}^{\prime})=\exp\left(\min\left\{0,W_{K}^{n}(x_{0:K}^{\prime})-W_{K}^{n}(x_{0:K})\right\}\right).$
When $K=0$, the accelerated swap coincides with the traditional PT swap in Equation˜2. However, an accelerated swap can obtain an acceptance of $1$ even if $\pi^{n-1}\neq\pi^{n}$ provided ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$. Therefore, we aim to choose the forward and backward accelerators to make the laws of simulated forward and backward paths as close to each other as possible. Theorem˜1 shows we can quantify this discrepancy in through the rejection rate, $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}):=\|{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}-{\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K}\|_{\mathrm{TV}}$ which by Pinsker inequality is controlled by the symmetric KL divergence between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$,
$r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})^{2}\leq\frac{1}{2}{\mathbb{P}}^{n-1}_{K}[-W^{n}_{K}]+\frac{1}{2}{\mathbb{Q}}^{n}_{K}[W^{n}_{K}]=:\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}).$ (4)
@@
Expectation and free energy estimators
As a by-product of Algorithm˜1, we obtain a consistent estimator $\hat{\pi}^{n}_{T}[f]=\frac{1}{T}\sum_{t=1}^{T}f(X^{n})$ for the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi^{n}$ by taking the Monte Carlo average over the $n$-th chain $X^{n}_{t}$. To also obtain free energy estimates, by averaging over $\accentset{\rightharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K})$ and $\accentset{\leftharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, the work done by the forward and backward accelerated paths generated during the $n$-th accelerated swap at time $t$ during the communication phase of Algorithm˜1, to obtain consistent estimators $\Delta\accentset{\rightharpoonup}{F}_{T}$ and $\Delta\accentset{\leftharpoonup}{F}_{T}$ for $\Delta F$ respectively as $T\to\infty$,
$\exp(-\Delta\accentset{\rightharpoonup}{F}_{T})\!:=\!\prod_{n=1}^{N}\frac{2}{T}\sum_{n\equiv t\mod 2}\exp\left(-\accentset{\rightharpoonup}{W}^{n}_{K,t}\right),~\exp(\Delta\accentset{\leftharpoonup}{F}_{T})\!:=\!\prod_{n=1}^{N}\frac{2}{T}\sum_{n\equiv t\mod 2}\exp\left(\accentset{\leftharpoonup}{W}^{n}_{K,t}\right).$
We then obtain a consistent estimator $\Delta\hat{F}_{T}=\frac{1}{2}(\Delta\accentset{\rightharpoonup}{F}_{T}+\Delta\accentset{\leftharpoonup}{F}_{T})$ for $\Delta F$ by averaging the forward and backward estimators. When we consider the classical PT with $K=0$, these estimators degrade back to free energy perturbation (FEP) (Zwanzig, 1954); when the path is deterministic (e.g., implemented through normalising flow), these estimators recover the target FEP (Jarzynski, 2002); and when the path is stochastic, the estimators can be viewed as a form of Jarzynski equality (Jarzynski, 1997) or escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). Additionally, following He et al. (2025a, Proposition 3.2), we can also apply Bennett acceptance ratio using the work for the paths (Bennett, 1976; Shirts et al., 2003; Hahn & Then, 2009; Minh & Chodera, 2009; Vaikuntanathan & Jarzynski, 2011) to achieve reduced variance without additional functional calls.
@@
The estimators $\hat{\pi}^{n}_{T}[f]$ and $\Delta\hat{F}_{T}$ a.s. converge to $\pi^{n}[f]$ and $\Delta F$ respectively as $T\to\infty$. Moreover, if ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$ for all $n$, then $\Delta\hat{F}_{T}\stackrel{{\scriptstyle a.s.}}{{=}}\Delta F$.

### v4

@@ Abstract
Markov Chain Monte Carlo (MCMC) algorithms are essential tools in computational statistics for sampling from unnormalised probability distributions, but can be fragile when targeting high-dimensional, multimodal, or complex target distributions. Parallel Tempering (PT) enhances MCMC’s sample efficiency through annealing and parallel computation, propagating samples from tractable reference distributions to intractable targets via state swapping across interpolating distributions. The effectiveness of PT is limited by the often minimal overlap between adjacent distributions in challenging problems, which requires increasing computational resources to compensate. We introduce a framework that accelerates PT by leveraging neural samplers—including normalising flows, diffusion models, and controlled diffusions—to reduce the required overlap. Our approach utilises neural samplers in parallel, circumventing the computational burden of neural samplers while preserving the asymptotic consistency of classical PT. We demonstrate theoretically and empirically on a variety of multimodal sampling problems that our method improves sample quality, reduces the computational cost compared to classical PT, and enables efficient free energy/normalising constant estimation.
@@ 1 Introduction
Sampling from a probability distribution $\pi(x)=\exp(-U(x))/Z$ defined over a state-space $\mathcal{X}$ with a tractable un-normalised density $\tilde{\pi}:\mathcal{X}\to\mathbb{R}$ and intractable normalising constant $Z=\int_{\mathcal{X}}\tilde{\pi}(x)\mathrm{d}x$ is a fundamental task in machine learning and natural sciences. Markov Chain Monte Carlo (MCMC) methods are usually employed for such purposes, constructing an ergodic Markov chain $(X_{t})_{t\in\mathbb{N}}$ using local moves leaving $\pi$ invariant. While MCMC algorithms are guaranteed to converge asymptotically, in practice, they struggle when the target is complex with multiple well-separated modes (Papamarkou et al., 2022; Hénin et al., 2022). To handle such cases, Parallel Tempering (PT) (Swendsen & Wang, 1986; Geyer, 1991; Hukushima & Nemoto, 1996) is a popular class of MCMC methods designed to improve the global mixing of locally efficient MCMC algorithms.
@@
PT works by considering an annealing path $\pi^{0},\pi^{1},\dots,\pi^{N}$ of distributions over $\mathcal{X}$ interpolating between a simple reference distribution $\pi^{0}=\eta$ (e.g., a Gaussian) and the target $\pi^{N}=\pi$. PT algorithms construct a Markov chain $\textbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$, targeting the joint distribution $\bm{\pi}=\pi^{0}\otimes\cdots\otimes\pi^{N}$. The PT chain $\mathbf{X}_{t}$ is constructed by alternating between (1) a local exploration phase where the $n$-th chain  [^1] Following the PT literature, we also refer to components of $\mathbf{X}_{t}$ as chains. of $\mathbf{X}_{t}$ is updated according to a $\pi^{n}$-invariant MCMC algorithm; and (2) a communication phase which proposes a sequence of swaps between neighbouring states accepted according to a Metropolis–Hastings correction ensuring invariance (see Figure˜1 (a)). Crucially, PT offsets the additional computation burden of simulating the extended $N$ chains through parallel computation, allowing for a similar effective computational cost as a single chain when implemented in a maximally parallelised manner.
@@
Typically, the chains $X_{t}^{n}$ mix faster when closer to the reference and struggle closer towards the target. Therefore, communication between the reference and target, facilitated through swaps, can induce rapid mixing between modes of the target component (Woodard et al., 2009; Surjanovic et al., 2024). While the importance of the swapping mechanism for PT has led to a literature dedicated to optimising communication between the reference and target (Syed et al., 2021; 2022; Surjanovic et al., 2022), such works still rely on the original swapping mechanism (Geyer, 1991).
@@
Figure 1: (Left) An illustration of the local exploration and communication step for PT vs APT. (Middle) 1,000 samples of a Gaussian mixture model target obtained using PT vs APT with a standard Gaussian reference. See Appendix 6.1 for more details. (Right) Round trips for PT and APT with $N=6$ chains over $T=100,000$ iterations of Algorithm 1.
@@
We answer positively this question by formalising and exploiting the framework introduced by Ballard & Jarzynski (2009; 2012) in physics for designing more flexible swap mechanisms, which we call Accelerated Parallel Tempering (APT). APT preserves PT’s asymptotic consistency and allows us to easily integrate normalising flows, diffusion models and stochastic control into existing PT implementations. Moreover, APT uses these neural samplers in a parallelised manner, mitigating their high computational burden. Empirically, APT outperforms PT by accelerating the communication between the reference and target states, even when controlling for the additional computation incurred by neural samplers.
A relevant prior work integrating PT with normalising flows is Invernizzi et al. (2022), which trains a flow to directly map configurations from the highest-temperature reference to the lowest-temperature target, effectively bypassing the intermediate annealing distributions. Abbott et al. (2024) also combines PT with normalising flows to accelerate sampling in lattice quantum chromodynamics; their approach is equivalent to the NF-APT scheme of Section˜5.1. Our framework is strictly more general, accommodating any learned or approximate transport map between neighbouring distributions.
@@ 2 Parallel Tempering
Let $\pi^{0},\pi^{1},\dots,\pi^{N}$ be an annealing path of probability distributions on $\mathcal{X}$ where $\pi^{0}=\eta$ is the reference and $\pi^{N}=\pi$ is the target. We assume the $n$-th annealing distribution admits density $\pi^{n}(x):=\tilde{\pi}^{n}(x)/Z_{n}$ with respect to a base measure $\mathrm{d}x$, where $\tilde{\pi}^{n}:\mathcal{X}\to\mathbb{R}$ is the un-normalised density which we can evaluate, with normalising constant $Z_{n}:=\int_{\mathcal{X}}\tilde{\pi}^{n}(x)\mathrm{d}x$. Our goal is to estimate $\pi[f]:=\int_{\mathcal{X}}f(x)\pi(\mathrm{d}x)$, the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi=\pi^{N}$, and the normalising constant $Z=Z_{N}$.
There is considerable flexibility in choosing annealing distributions provided $\pi^{0}=\eta$ with $Z_{0}=1$ and $\pi^{N}=\pi$ with $Z_{N}=Z$. Without loss of generality, we can assume $\pi^{n}=\pi^{\beta_{n}}$ where for $\beta\in[0,1]$, $\pi^{\beta}$ continuously interpolates between the reference and target as $\beta$ increases from 0 to $1$, according to some annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$. A common choice is the geometric path, $\pi^{\beta}(x)\propto\eta(x)^{1-\beta}\pi(x)^{\beta}$, which linearly interpolates between reference and target in log-space. See Masrani et al. (2021); Syed et al. (2021); Máté & Fleuret (2023); York (2023) for alternative non-geometric annealing paths.
@@ 2.1 Non-Reversible Parallel Tempering
The PT algorithm constructs a Markov chain $\mathbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$ invariant to the joint distribution $\bm{\pi}:=\pi^{0}\otimes\cdots\otimes\pi^{N}$. We construct $\mathbf{X}_{t}$ from $\mathbf{X}_{t-1}$ by doing (1) a local exploration step followed by (2) a communication step seen in the top of Figure˜1 (a). For $n=0,\dots,N$, the $n$-th local exploration move (1) updates the $n$-th component of $\mathbf{X}_{t-1}$ using a $\pi^{n}$-invariant Markov kernel $K^{n}(x,\mathrm{d}x^{\prime})$ on $\mathcal{X}$ corresponding to an MCMC move targeting $\pi^{n}$,
@@
$\alpha^{n}(x,x^{\prime}):=\min\left\{1,\frac{w^{n}(x^{\prime})}{w^{n}(x)}\right\}.$ (1)
Where $w^{n}:\mathcal{X}\to\mathbb{R}$ is the incremental weight equal to the un-normalised Radon-Nikodym derivative:
$w^{n}(x):=\frac{Z_{n}}{Z_{n-1}}\frac{\mathrm{d}\pi^{n}}{\mathrm{d}\pi^{n-1}}(x)=\frac{\tilde{\pi}^{n}(x)}{\tilde{\pi}^{n-1}(x)}.$ (2)
In practice, it is advantageous to use a non-reversible communication (Okabe et al., 2001; Syed et al., 2022), where the $n$-th swap move is proposed only at iterations with matching parity: $n\equiv t\mod 2$. Both local exploration and communication steps can be done in parallel, allowing distributed implementations to leverage parallel computation to accelerate sampling (Surjanovic et al., 2023).
Round trips.
While the effective sample size (ESS) of samples generated by a Markov chain is the gold standard for evaluating the performance of MCMC algorithms, in our setting, we are mainly interested in improving the swap kernel within PT. As ESS measures the intertwined performance of the local exploration and swap kernels, we instead evaluate the performance of PT and APT in terms of the communication between reference and target, in order to disentangle the influence of the swap kernels from the local exploration. This is empirically measured by counting the total number of round trips $R_{T}$ which tracks the number of independent reference samples transported to the target after $T$ iterations of PT (Katzgraber et al., 2006; Lingenheil et al., 2009); see Appendix A for a formal definition. In particular, the mixing time of PT is related to the time it takes for a round trip to occur and the total number of round trips is a measure of the particle diversity generated by PT and strongly correlates with the ESS (Surjanovic et al., 2024).
@@
12:   if $\log U<\log\accentset{\rightharpoonup}{w}^{n}_{K,t}-\log\accentset{\leftharpoonup}{w}^{n}_{K,t}$ then $\triangleright$ Accelerated swap move
13:     $X^{n-1}_{t},X^{n}_{t}\leftarrow\accentset{\leftharpoonup}{X}^{n}_{t,0},\accentset{\rightharpoonup}{X}^{n-1}_{t,K}$
@@ 3 Accelerated Parallel Tempering
A limitation of PT is its inflexible swap move, which only proposes directly exchangeable samples between distributions, considering just the relative change in likelihood $\pi^{n-1}$ and $\pi^{n}$. Low acceptance probability occurs when these distributions have minimal overlap. Addressing this requires increasing the number of parallel chains $N$ which may not always be possible. We propose Accelerated Parallel Tempering (APT), expanding the framework developed in Ballard & Jarzynski (2009; 2012), to improve distributional overlap and accelerate communication for a fixed number of chains.
@@ 3.1 Forward and Backward Accelerators
For $n=1,\dots,N$, let $P_{k}^{n-1}(x_{k-1},\mathrm{d}x_{k})$ and $Q_{k-1}^{n}(x_{k},\mathrm{d}x_{k-1})$, $k=1,\dots,K$, be two families of $K$ transition kernels which we call forward accelerators and backward accelerators. They induce ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$, two time-inhomogeneous Markov processes obtained by propagating $\pi^{n-1}$ forward in time and $\pi^{n}$ backward in time respectively using the forward and the backward accelerators:
@@
We assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ are mutually absolutely continuous and we can evaluate $w^{n}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$, the incremental weights between the forward and the backward paths, extending $w^{n}$ in Equation˜2,
$w^{n}_{K}(x_{0:K}):=\frac{Z_{n}}{Z_{n-1}}\frac{\mathrm{d}{{\mathbb{Q}}}^{n}_{K}}{\mathrm{d}{{\mathbb{P}}}^{n-1}_{K}}(x_{0:K}).$ (3)
@@ 3.2 Non-Reversible Accelerated Parallel Tempering
We construct the Markov chain $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ for $t=1,\dots,T$ using the same local exploration and non-reversible communication as classical PT, but we use the accelerated PT swap move as described below and summarized in Algorithm˜1. Given PT state $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})\in\mathcal{X}^{N+1}$ after the local exploration move, we define the $n$-th accelerated swap move as follows: generate paths $\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K}$, and $\accentset{\leftharpoonup}{X}^{n}_{t,0:K}$ obtained by propagating $X^{n-1}_{t}$ and $X^{n}_{t}$ forward and backward in time using the forward and backward transitions respectively,
@@
$\alpha_{K}^{n}(x_{0:K},x_{0:K}^{\prime})=\min\left\{1,\frac{w^{n}_{K}(x^{\prime}_{0:K})}{w^{n}_{K}(x_{0:K})}\right\}.$
When $K=0$, the accelerated swap coincides with the traditional PT swap in Equation˜1. However, an accelerated swap can obtain an acceptance of $1$ even if $\pi^{n-1}\neq\pi^{n}$ provided ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$. Therefore, we aim to choose the forward and backward accelerators to make the laws of simulated forward and backward paths as close to each other as possible. Theorem˜1 shows we can quantify this discrepancy through the (theoretical) rejection rate, $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}):=\|{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}-{\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K}\|_{\mathrm{TV}}$ which by Pinsker inequality is controlled by the symmetric KL divergence between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ (for the proof, see Appendix B.1):
$r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})^{2}\leq\frac{1}{2}{\mathbb{P}}^{n-1}_{K}[-\log w^{n}_{K}]+\frac{1}{2}{\mathbb{Q}}^{n}_{K}[\log w^{n}_{K}]=:\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}).$ (4)
@@
Expectation and free energy estimators.
As a by-product of Algorithm˜1, we obtain a consistent estimator $\hat{\pi}^{n}_{T}[f]=\frac{1}{T}\sum_{t=1}^{T}f(X^{n}_{t})$ for the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi^{n}$ by taking the Monte Carlo average over the $n$-th chain $X^{n}_{t}$. To also obtain free energy estimates, we can average over $\accentset{\rightharpoonup}{w}^{n}_{K,t}:=w^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K})$ and $\accentset{\leftharpoonup}{w}^{n}_{K,t}:=w^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, the weights of the forward and backward accelerated paths generated during the $n$-th accelerated swap at time $t$ during the communication phase of Algorithm˜1, to obtain consistent estimators $\accentset{\rightharpoonup}{Z}_{T}$ and $\accentset{\leftharpoonup}{Z}_{T}$ for $Z$ respectively as $T\to\infty$,
$\accentset{\rightharpoonup}{Z}_{T}:=\prod_{n=1}^{N}\frac{2}{T}\sum_{n\equiv t\mod 2}\accentset{\rightharpoonup}{w}^{n}_{K,t},\quad\frac{1}{\accentset{\leftharpoonup}{Z}_{T}}:=\!\prod_{n=1}^{N}\frac{2}{T}\sum_{n\equiv t\mod 2}\frac{1}{\accentset{\leftharpoonup}{w}^{n}_{K,t}}.$
We then obtain a consistent estimator $\hat{Z}_{T}=(\accentset{\rightharpoonup}{Z}_{T}\accentset{\leftharpoonup}{Z}_{T})^{1/2}$ for $Z$ by taking the geometric mean of the forward and backward estimators. Proposition 1 provides the formal statement of these results; for the proof, see Appendix B.1.3.
@@
The estimators $\hat{\pi}^{n}_{T}[f]$ and $\hat{Z}_{T}$ a.s. converge to $\pi^{n}[f]$ and $Z$ respectively as $T\to\infty$. Moreover, if ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$ for all $n$, then $\hat{Z}_{T}\stackrel{{\scriptstyle a.s.}}{{=}}Z$.
When we consider the classical PT with $K=0$, these estimators degrade back to free energy perturbation (FEP) (Zwanzig, 1954); when the path is deterministic (e.g., implemented through normalising flow), these estimators recover the target FEP (Jarzynski, 2002); and when the path is stochastic, the estimators can be viewed as a form of Jarzynski equality (Jarzynski, 1997) or escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). Additionally, following He et al. (2025a, Proposition 3.2), we can also apply the Bennett acceptance ratio using the weight for the paths (Bennett, 1976; Shirts et al., 2003; Hahn & Then, 2009; Minh & Chodera, 2009; Vaikuntanathan & Jarzynski, 2011) to achieve reduced variance without additional functional calls.